# Personalized Heart Disease Detection via ECG Digital Twin Generation

### v1

@@ Abstract
Heart diseases rank among the leading causes of global mortality, demonstrating a crucial need for early diagnosis and intervention. Most traditional electrocardiogram (ECG) based automated diagnosis methods are trained at population level, neglecting the customization of personalized ECGs to enhance individual healthcare management. A potential solution to address this limitation is to employ digital twins to simulate symptoms of diseases in real patients. In this paper, we present an innovative prospective learning approach for personalized heart disease detection, which generates digital twins of healthy individuals’ anomalous ECGs and enhances the model sensitivity to the personalized symptoms. In our approach, a vector quantized feature separator is proposed to locate and isolate the disease symptom and normal segments in ECG signals with ECG report guidance. Thus, the ECG digital twins can simulate specific heart diseases used to train a personalized heart disease detection model. Experiments demonstrate that our approach not only excels in generating high-fidelity ECG signals but also improves personalized heart disease detection. Moreover, our approach ensures robust privacy protection, safeguarding patient data in model development. The code can be found at https://github.com/huyjj/LAVQ-Editor
1 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_introduction
2 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_related
3 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_method_1
4 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_exp
5 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_conclution
@@ Acknowledgments
This research was partially supported by National Natural Science Foundation of China under grants No. 62176231, No. 62106218, No. 82202984, No. 92259202 and No. 62132017, Zhejiang Key R&D Program of China under grant No. 2023C03053.
Appendix A: Ablation Study of threshold $l$
@@
Appendix B: Extra Visualization Results
Referring to Figures  2,  4, and  4, we display and intuitively inspect the generated ECGs with heart diseases: Hypertrophy, Myocardial Infarction, and Conduction Disturbance, by various models. Our findings reveal:
@@
(iii) Our method successfully generates correct heartbeat features and markedly reduces noise, boosting ECG realism. Furthermore, our approach distinctly reveals clear pathological signs. As indicated by orange rectangles in Figure  2, we observe T-wave flattening in lead I, indicative of potential hypertrophic cardiomyopathy. Figure  4 display pronounced QRS abnormalities in lead V3, where the T wave is notably higher than the R and S waves. Figure  4 shows instances of the P wave failing to progress to the ventricle, potentially indicative of ventricular or junctional escape rhythms.

### v2

@@ Abstract
Heart diseases rank among the leading causes of global mortality, demonstrating a crucial need for early diagnosis and intervention. Most traditional electrocardiogram (ECG) based automated diagnosis methods are trained at population level, neglecting the customization of personalized ECGs to enhance individual healthcare management. A potential solution to address this limitation is to employ digital twins to simulate symptoms of diseases in real patients. In this paper, we present an innovative prospective learning approach for personalized heart disease detection, which generates digital twins of healthy individuals’ anomalous ECGs and enhances the model sensitivity to the personalized symptoms. In our approach, a vector quantized feature separator is proposed to locate and isolate the disease symptom and normal segments in ECG signals with ECG report guidance. Thus, the ECG digital twins can simulate specific heart diseases used to train a personalized heart disease detection model. Experiments demonstrate that our approach not only excels in generating high-fidelity ECG signals but also improves personalized heart disease detection. Moreover, our approach ensures robust privacy protection, safeguarding patient data in model development. The code can be found at https://github.com/huyjj/LAVQ-Editor.
@@ Acknowledgments
This research was supported by Zhejiang Provincial Natural Science Foundation of China (LY23F020019), National Natural Science Foundation of China (62106218 & 82200558), Excellent Young Scientists Fund (82122036), Basic and Applied Basic Research Foundation of Guangdong Province (2022A1515110722), and Guangdong Provincial Medical Science & Technology Research Fund Project (A2023011).
Appendix A: Pseudocode of Training
  Initialize VQ-Separator, Generator $G$, and Discriminator $D$
  for each training epoch do
     for each batch in the dataset do
        $X_{ecg}\leftarrow$ pre-diagnosis ECG, $X_{text}\leftarrow$ the report of ECG
        $X_{ecg}^{t}\leftarrow$ reference ECG, $X_{text}^{t}\leftarrow$ the report of reference ECG
        Optimizer Discriminator $D$:
        $f_{d},f_{p}\leftarrow\text{VQ-Separator}({X}_{ecg},X_{text})$
        $f_{d}^{t},f_{p}^{t}\leftarrow\text{VQ-Separator}({X}_{ecg}^{t},X_{text}^{t})$
        $\hat{X_{ecg}^{t}}\leftarrow G_{dec}(f_{p},f_{d}^{t})$
        $\hat{X}_{ecg}^{rec}\leftarrow G_{dec}(f_{p},f_{d})$
        $f_{d}^{G},f_{p}^{G}\leftarrow\text{VQ-Separator}(\hat{X}_{ecg}^{t},X_{text}^{t})$
        $\hat{f_{d}},D_{real}\leftarrow D(X_{ecg})$
        $\hat{f_{d}^{rec}},D_{fake}\leftarrow D(\hat{X}_{ecg}^{rec})$
        Calculate losses:
        $L_{adv}^{D}\leftarrow\log(D_{real})+\log(1-D_{fake})$
        $L_{sim}^{D}\leftarrow\mathcal{L}_{sim}^{D}=1-\frac{f_{d}^{D}\cdot f_{d}}{\parallel f_{d}^{D}\parallel\parallel f_{d}\parallel}$
        Update $D$ by minimizing $L_{sim}^{D}+L_{adv}^{D}$
        Optimizer Generator $G$:
        $f_{d},f_{p},L_{vq}\leftarrow\text{VQ-Separator}({X}_{ecg},X_{text})$
        $f_{d}^{t},f_{p}^{t}\leftarrow\text{VQ-Separator}({X}_{ecg}^{t},X_{text}^{t})$
        $\hat{X_{ecg}^{t}}\leftarrow G_{dec}(f_{p},f_{d}^{t})$
        $\hat{X_{ecg}^{rec}}\leftarrow G_{dec}(f_{p},f_{d})$
        $f_{d}^{G},f_{p}^{G}\leftarrow\text{VQ-Separator}(\hat{X}_{ecg}^{t},X_{text}^{t})$
        $\hat{f_{d}^{rec}},D_{fake}\leftarrow D(\hat{X}_{ecg}^{rec})$
        Calculate losses:
        $L_{recon}\leftarrow||X_{ecg}-\hat{X}_{ecg}^{rec}||^{2}$
        $L_{adv}\leftarrow\log(1-D_{fake})$
        Update $G$ by minimizing $L_{recon}+L_{vq}+L_{adv}^{G}+L_{sim}^{G}$
     end for
     Evaluate model performance on validation set
  end for
Algorithm 1 Training Process of the LAVQ-Editor
Appendix B: Ablation Study of threshold $l$
@@
Appendix C: Extra Visualization Results
Referring to Figures  3,  5, and  5, we display and intuitively inspect the generated ECGs with heart diseases: Hypertrophy, Myocardial Infarction, and Conduction Disturbance, by various models. Our findings reveal:
@@
(iii) Our method successfully generates correct heartbeat features and markedly reduces noise, boosting ECG realism. Furthermore, our approach distinctly reveals clear pathological signs. As indicated by orange rectangles in Figure  3, we observe T-wave flattening in lead I, indicative of potential hypertrophic cardiomyopathy. Figure  5 display pronounced QRS abnormalities in lead V3, where the T wave is notably higher than the R and S waves. Figure  5 shows instances of the P wave failing to progress to the ventricle, potentially indicative of ventricular or junctional escape rhythms.
For timing problems, as shown in Figure 2, our model can scale the signals based on the statistical RR interval length after morphological editing.
Figure 2: Example of ECG described as “sinus tachycardia apart from rate, normal ecg.” generated from the real normal ECG.